OpenStudy (anonymous):

Complete the equation of the graphed linear function in point-slope form. y – 8 = (x – ) The graph of the function goes through point 2, 8 and point 6, 0.

OpenStudy (anonymous):

@jigglypuff314 can you help me though this one to?:)

OpenStudy (igreen):

Find the slope. (2, 8) and (6, 0) \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{0-8}{6-2}\) Simplify

OpenStudy (anonymous):

ok 1 sec

OpenStudy (anonymous):

-8/4=-2?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yep, you got it that's our slope, -2. Now plug that into point-slope form along with (2, 8). \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value of the point, \(x_1\) is the x-value of the point, and \(m\) is the slope. Can you do that?

OpenStudy (anonymous):

Y-8=-2x-2?

OpenStudy (anonymous):

@iGreen Y-8=-2x-2?

OpenStudy (igreen):

y - 8 = -2(x - 2) Yes, that's correct.

OpenStudy (igreen):

\(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\) You can give medals by clicking 'Best Response'. @alyssa831